Give first the step you will use to separate the variable and then solve the equation:
p/7= 4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Now we have to multiply both sides of the equation by 7,
Then we get,
= p/7 × 7= 4 × 7
= p = 28

Thanks